### v1

# OrbitZoo: Multi-Agent Reinforcement Learning Environment for Orbital Dynamics

Abstract
The increasing number of satellites and orbital debris has made space congestion a critical issue, threatening satellite safety and sustainability. Challenges such as collision avoidance, station-keeping, and orbital maneuvering require advanced techniques to handle dynamic uncertainties and multi-agent interactions. Reinforcement learning (RL) has shown promise in this domain, enabling adaptive, autonomous policies for space operations; however, many existing RL frameworks rely on custom-built environments developed from scratch, which often use simplified models and require significant time to implement and validate the orbital dynamics, limiting their ability to fully capture real-world complexities. To address this, we introduce OrbitZoo, a versatile multi-agent RL environment built on a high-fidelity industry standard library, that enables realistic data generation, supports scenarios like collision avoidance and cooperative maneuvers, and ensures robust and accurate orbital dynamics. The environment is validated against a real satellite constellation, Starlink, achieving a Mean Absolute Percentage Error (MAPE) of 0.16% compared to real-world data. This validation ensures reliability for generating high-fidelity simulations and enabling autonomous and independent satellite operations.
March 2025
1 Introduction
Since the dawn of the space age in 1957, humanity has successfully launched approximately 20,000 satellites into Earth’s orbit [18], of which only about 50% remain operational. These satellites are crucial for our daily life, providing critical services such as global communication, navigation, weather forecasting, Earth observation, and scientific research. However, these advances come with many problems.
Earth’s orbital environment hosts an estimated 140 million debris objects, with approximately 1 million of these being larger than 1 cm – large and fast enough to cause catastrophic damage upon impact [17]. Collisions with space debris generate more debris, leading to further collisions. This chain-reaction known as the Kessler syndrome, results in an exponential increase in debris, threatening the long-term sustainability of Earth’s orbits [27]. If no further measures are taken to address this issue, Earth’s orbits could become unusable.
More recently, the congestion in Lower Earth Orbit (LEO) is likely to undergo a major change of scale for mega-constellations of telecommunication satellites, leveraging tens of thousands of satellites [9]. The increasing density of satellites and debris, particularly in LEO [11], presents formidable challenges for Space Traffic Management (STM), related to the accurate monitoring and tracking of space objects and development of adequate decision-support frameworks [35], compounded by issues of data scarcity and uncertainty. Despite the growing availability of orbital data and tools, current solutions for satellite maneuvers – ranging from orbital transfers to collision avoidance – remain heavily reliant on manual processes. Human experts must navigate an ever-increasing complexity of scenarios, making critical decisions under time constraints with incomplete or inaccurate information [21]. As the volume of satellites and orbital debris continues rising, these traditional methods are quickly becoming unsustainable, which demands the development of new strategies to minimize our impact, such as debris removal [7] and faster, more capable, and autonomous intelligent systems for decision-making. Adaptive control systems [4] provide robustness to uncertainty but struggle with complex, dynamic, and nonlinear space systems. RL, on the other hand, excels in real-time adaptation to such environments [49].
We explore how the application of RL to orbital dynamics is shaped by two primary factors and suffers from a lack of standardization. First, various tools are available to generate data, influencing the balance between simplicity and realism in the environment. Second, frameworks for developing RL-based missions within these environments are often built from scratch, calling for additional validation of the dynamics – a particularly challenging task in the field of orbital mechanics. To enable the effective application of RL to satellite maneuvering, we created OrbitZoo, an environment designed for both high-fidelity orbital data generation and RL development. The contributions of this paper are structured into three main categories:
  • Data Generation: Built on Python and with a robust space dynamics library on its background, OrbitZoo generates high-fidelity orbital data by incorporating realistic forces and perturbations, providing accurate datasets essential for machine learning and strategy validation;
  • Reinforcement Learning: OrbitZoo is standardized for RL research, leveraging the PettingZoo library [48] to support multi-agent reinforcement learning (MARL) with a Partially Observable Markov Decision Process (POMDP) structure. This integration enables the development, training, and benchmarking of intelligent satellite maneuvering strategies for single and multi-agent missions in cooperative, competitive, or mixed scenarios;
  • Customizable Framework with Visualization: OrbitZoo’s modular design allows users to define scenarios, incorporate custom models, and adapt the environment to specific needs, with clear separation of abstraction levels. It also features an interactive visual component, making it versatile for a wide range of applications while providing an accessible way to understand orbital behaviors and decision-making processes.
2 Related Work
Existing efforts in this domain can be broadly categorized into four areas: high-fidelity tools for generating and simulating orbital dynamics data, reinforcement learning (RL) for orbital dynamics, multi-agent RL (MARL) environments, and efforts to bridge the reality gap.
Data Generation and Simulation Tools.
Orekit [33] is one of the most comprehensive open-source libraries for astrodynamics and orbital mechanics, offering advanced capabilities for precise orbit determination [34, 38, 14], propagation and associated uncertainties [2], attitude determination, and trajectory analysis [16]. Its modular, Java-based architecture allows users to model systems ranging from simple Newtonian attraction to highly complex scenarios incorporating detailed gravity fields and perturbations, such as atmospheric drag, solar radiation pressure, and third-body effects. While versatile, Orekit’s steep learning curve and reliance on Java can pose challenges for users unfamiliar with its technicalities.
Poliastro [39], a Python-native library, offers a more accessible alternative, featuring tools for orbit propagation and transfer planning. Although it lacks the high-fidelity modeling of Orekit, Poliastro integrates with the Cesium library [15] for accurate 3D geospatial visualization, making it suitable for simpler scenarios or users with limited expertise in orbital mechanics.
Systems Tool Kit (STK) [5], the leading commercial orbital simulation software, offers extensive capabilities for mission planning and operational analysis. Despite its advanced visualization features and high precision, STK’s substantial cost restricts its accessibility to well-funded organizations.
Reinforcement Learning for Orbital Dynamics.
Reinforcement learning has been widely applied to satellite maneuvering, leveraging custom-built environments and advanced dynamics models. Researchers have frequently used the Circular Restricted Three-Body Problem (CR3BP) model [36, 19, 46, 29] for simplified simulations, though it often fails to capture real-world dynamics and thrust perturbations, posing challenges for non-experts in orbital dynamics aiming to extend RL to realistic scenarios. For Low-Earth Orbit (LEO) station-keeping, [6] developed an RL framework using PPO with a 4th-order Yoshida integrator that accounted for gravitational forces and atmospheric drag. [13] explored orbital maneuvers in geostationary orbit (GEO), using A2C for a perigee raise. However, both studies relied on simplified dynamics, limiting their applicability to more complex scenarios. Orekit has also been central to RL environments tackling complex dynamics. For instance, [28] employed a DDPG-based framework for low-thrust transfers but lacked adaptability to multi-agent scenarios and dynamic uncertainties. Collision Avoidance Maneuvers (CAMs) have been another key focus. [26] introduced ColAvGym, a single-agent environment for CAM in LEO, leveraging Orekit for dynamics and PPO for learning. Using real Conjunction Data Messages (CDMs), the environment retroactively reconstructed satellite and debris orbits for training. Alternatively, [45] and [12] relied on synthetic methods to generate debris scenarios, using environments built on OpenAI’s Gym with Orekit and a Simplified General Perturbations model (SGP4), respectively. These studies explored various observation and reward designs, including temporal features captured through recurrent networks like DRQN, to handle the problem as a Partially Observable Markov Decision Process (POMDP).
Multi-Agent Reinforcement Learning Environments
MARL has recently gained traction in orbital dynamics, with applications in satellite coordination and task assignment. [23] introduced the RL-Enabled Distributed Assignment (REDA) algorithm, using Poliastro and DQN for task allocation across satellite constellations. However, REDA lacked integration with realistic orbital data. Similarly, [53] applied Multi-Agent PPO (MAPPO) for multi-satellite observation planning, combining STK and MATLAB for custom simulation frameworks but without leveraging industry-standard astrodynamics libraries like Orekit.
Bridging the Reality Gap
A critical challenge in applying RL to orbital dynamics is bridging the Reality Gap between simulation and real-world operations. This includes incorporating high-fidelity dynamics, handling uncertainties, and validating simulation results against real-world data. For example, [26] developed ColAvGym, which used real Conjunction Data Messages (CDMs) to train agents for collision avoidance. [45] and [12] relied on synthetic data and simplified dynamics for training, but their environments did not fully capture the complexities of realistic orbital interactions.
Table 1: Comparison of RL environments for orbital dynamics.
Work Multi-Agent RL Industry-Standard Simulator High-Fidelity Dynamics Continuous Control Realistic Bodies and Thrust Interactive Visualization Publicly Available
Kolosa (2019)
Miller (2019)
Armando (2020) ✓(drag only)
Federici (2021)
Sullivan (2021)
Casas (2022)
Bourriez (2023)
LaFarge (2023) ✓(third bodies only)
Zhang (2023)
Solomon (2024)
Holder (2024)
Kazemi (2024)
OrbitZoo (ours)
As shown in Tab. 1, OrbitZoo stands out with multi-agent RL support, high-fidelity orbital dynamics, and integration with Orekit, an industry-standard simulator. Unlike prior works with simplified models, OrbitZoo incorporates perturbative forces, supports Cartesian and equinoctial representations, and offers realistic body dynamics, thrust modeling, and interactive visualization. Its public availability further establishes it as a comprehensive framework for advancing learning-based approaches in space operations.
3 Background: Challenges in Multi-Agent RL for Orbital Dynamics
Reinforcement Learning (RL) provides a framework for decision-making under uncertainty, where agents learn through trial and error to maximize cumulative rewards [47, 43]. In the context of multi-agent RL (MARL), multiple agents operate in a shared environment, each pursuing individual or collective goals. MARL presents unique challenges, particularly in orbital dynamics, where environments are partially observable, highly dynamic, and governed by complex physical interactions.
3.1 Theoretical Foundations of MARL
MARL builds upon Markov Decision Processes (MDPs) [47, 10] and Partially Observable Markov Decision Processes (POMDPs), commonly used to model scenarios where agents cannot access the full environment state [25]. A POMDP is defined by the tuple $\langle S,A,P,R,O,\Omega\rangle$. At each timestep $t$, an agent observes $o_{t}\in O$ based on the current state $s_{t}\in S$, selects an action $a_{t}\in A$, and transitions to a new state $s_{t+1}$ with probability $P(s_{t+1}|s_{t},a_{t})$, receiving a reward $R(s_{t},a_{t})$. The goal is to learn a policy $\pi(a_{t}|o_{t})$ that maximizes expected returns over time.
Deep RL algorithms such as DDPG [31] and PPO [42] have been instrumental in extending traditional RL to continuous control problems, albeit the topic had been approached before [44]. In MARL, the interaction between agents is often modeled as a multi-agent POMDP (MA-POMDP), where each agent $i$ observes $o^{i}_{t}$, performs an action $a^{i}_{t}$, and receives a reward $r^{i}_{t}$. Proximal Policy Optimization (PPO) and variants, such as Multi-Agent PPO (MAPPO) [52], are well-suited for these scenarios, allowing a stable decentralized or centralized training of agents in shared environments. This is particularly important for missions involving constellations of satellites, where each satellite may need to coordinate, compete, or cooperate with others to achieve shared objectives in complex – and potentially noisy – orbital dynamics.
3.2 Challenges in Bridging the Reality Gap
Simulating orbital dynamics introduces a significant reality gap due to the need for high-fidelity physics modeling and integration of real-world uncertainties. This section outlines key challenges encountered when applying RL to orbital dynamics. Basic concepts related to orbital mechanics are explored in 0.A.
Complex Orbital Dynamics
Orbital dynamics are governed by Newton’s laws of motion and universal gravitation, requiring accurate modeling of perturbative forces, including atmospheric drag, solar radiation pressure, and third-body gravitational effects. Simplified propagators, such as the Simplified General Perturbations (SGP) model, are computationally efficient but fail to capture the precision required for RL tasks. As provided by Orekit [33], numerical propagators offer higher fidelity by incorporating complex perturbations, enabling realistic simulations for long-term trajectory predictions and mission-critical maneuvers.
Coordinate Systems and State Representations
Orbital states are traditionally represented using Keplerian or Cartesian coordinates. However, these systems suffer from singularities in special cases, such as circular or equatorial orbits. Equinoctial elements are preferred in RL applications as they avoid such singularities while maintaining the advantages of traditional representations.
Thrust Modeling and Control
Thrust actions are typically modeled in local reference frames such as RSW (Radial, Along-track, and Cross-track). Polar parametrization offers a physically realistic and constrained action space, where the thrust is defined in magnitude, deviation angle, and azimuthal angle, providing greater flexibility for modeling satellite maneuvering capabilities. The transformation between polar and Cartesian thrust vectors ensures compatibility with existing numerical propagators.
Ephemeris Data and Validation
Ephemerides provide time-stamped predictions of orbital positions and velocities, enabling the validation of simulated trajectories. For example, we show that Orekit’s numerical propagator performs well when validated against ephemeris data from Starlink satellites. By optimizing physical parameters, such as drag and reflection coefficients, the propagator closely approximates real-world satellite trajectories, helping to bridge the reality gap in RL applications.
3.3 Exploration in High-Dimensional Action Spaces
Exploration in continuous action spaces, such as those encountered in orbital dynamics, presents additional challenges. Traditional RL algorithms like DDPG rely on deterministic policies and Gaussian noise for exploration, which may be suboptimal in complex environments. Algorithms like PPO address this limitation by outputting stochastic policies, allowing agents to explore action spaces and learn optimal maneuvers adaptively.
3.4 Multi-Agent Coordination and Scalability
In MARL for orbital dynamics, agents often face partially observable environments with decentralized policies. Challenges include coordinating tasks among agents, avoiding interference, and scaling solutions to larger constellations. Recent works address these challenges by using C++ engines [54], or CPU/GPU parallelization for faster agent training [40] and step/propagation time [30]. The EPyMARL library [37] builds on PyMARL to enable flexible multi-agent configurations, allowing satellite collaboration. However, MARLlib [24], which integrates with PettingZoo environments like OrbitZoo, may be more suitable, as it offers more MARL algorithms and neural network architectures (e.g., recurrent and convolutional).
Centralized critics and techniques like generalized advantage estimation (GAE) [41] improve training stability and efficiency for such environments, enabling effective multi-agent coordination.
3.5 Realistic Simulation Environments
Realistic simulation environments, such as OrbitZoo, must integrate high-fidelity physics models, flexible state representations, and scalable multi-agent frameworks. They should also provide interactive visualization tools and be publicly accessible to enable reproducibility and extensibility. By addressing the limitations of existing environments (e.g., lack of multi-agent support or simplified dynamics), OrbitZoo aims to advance learning-based approaches in space operations.
4 OrbitZoo: A Framework for Multi-Agent RL in Orbital Dynamics
Figure 1: High-level representation of OrbitZoo’s architecture. A JSON with information regarding a specific system must be provided to the environment (OrbitZoo). The environment then acts as the interface for generating data, developing single and multi-agent RL missions, or simply analyzing orbital dynamics.
OrbitZoo is a flexible and modular environment designed to address the challenges of applying multi-agent reinforcement learning (MARL) to high-fidelity orbital dynamics. It overcomes key limitations in existing tools, such as a lack of configurability, restricted support for realistic perturbative forces, and limited multi-agent capabilities. By integrating with Orekit, an industry-standard library for orbital mechanics, OrbitZoo ensures accuracy in physical modeling while remaining extensible for a wide range of RL missions. Its architecture, illustrated in Figure 1, provides modular components for data generation, mission design, and visualization, enabling researchers to systematically address the reality gap.
4.1 Architecture and Design
OrbitZoo’s architecture is designed to be modular and extensible. Figure 1 shows the primary modules, where gray boxes represent developed classes and other colors indicate integrated external components. This modular design allows users to implement and customize each component independently, ensuring compatibility with diverse RL algorithms and experimental setups.
4.2 Core Modules
Body: Modular Propagation of Orbital Dynamics
The Body class is the foundation of OrbitZoo and represents physical entities in the environment. Each body instance interfaces directly with Orekit’s numerical propagator, enabling high-fidelity simulations of orbital dynamics. The Body class incorporates: a unique identifier (alphanumeric value), a dry mass (in kg), a radius (in m), an expected Cartesian position and velocity $\mu=(x,y,z,\dot{x},\dot{y},\dot{z})$, and uncertainties associated with each of these elements $(\sigma_{x},\sigma_{y},\sigma_{z},\sigma_{\dot{x}},\sigma_{\dot{y}},\sigma_{\dot{z}})$, which are internally used to construct the covariance matrix $\Sigma$. Optionally, the initial date can be provided, enabling precise simulation of forces acting on the satellite at a specific moment. When resetting the environment, the actual position and velocity of the body are sampled from a multivariate normal distribution $\mathcal{N}(\mu,\Sigma)$. There are also static methods that provide information related to two given bodies, such as the distance between them, time of closest approach or probability of collision.
Satellite: Extending Bodies with Thrust Capabilities
The Satellite class extends the Body class to incorporate propulsion systems and agent parameters, essential for RL tasks requiring control and maneuverability. Satellites are configured with: (1) thrust parameters, including maximum thrust $(T_{\text{max}})$, deviation angle $(\theta_{\text{max}})$, and azimuthal angle $(\phi_{\text{max}})$ for realistic thrust modeling; (2) initial fuel mass and thrust-specific impulse for long-duration missions; (3) polar thrust parametrization, allowing satellites to apply thrust in any direction within physical constraints. This design enables RL agents to explore and optimize complex multi-dimensional action spaces; and (4) agent parameters. The agent parameters are arbitrary and serve as a standardization, depending on the specific requirements of the implemented algorithm for initialization.
Interface: Interactive Visualization
The Interface class provides interactive 3D visualization of the orbital environment, making OrbitZoo particularly useful for debugging, analysis, and presentation. It supports: (1) customizable visual components, such as equatorial grids, velocity and thrust vectors, and satellite trails; (2) real-time updates of system states, including timestamps, body names, and orbital parameters; and (3) flexible camera perspectives for inspecting multi-agent interactions and orbital trajectories.
(a) A system without bodies but with 3 different orbits.
(b) A system propagating several Starlink satellites.
(c) A single-agent mission of a satellite following a leader.
(d) A multi-agent mission of a constellation in GEO.
Figure 2: Frames of different systems on OrbitZoo’s interface.
Figure 2 demonstrates the visualization capabilities, ranging from single-agent missions to multi-agent constellations.
Additionally, we demonstrate this tool through four videos that highlight its capabilities in various scenarios. The first video offers an overview of the user interface, followed by demonstrations of the Hohmann maneuver mission, the GEO constellation mission, and a chase-target mission. These videos can be accessed through the following links: Interface Video; Hohmann Maneuver Mission; GEO Constellation Mission; Chase Target Mission.
The interface is built upon play3d, a library that leverages 2D perspective projections to create interactive 3D environments. Play3d is supported by the well-known 2D library pygame, which serves as its foundation.
Environment: High-Level Interaction
The Environment (class OrbitZoo) serves as the primary interface for users. It integrates the above modules to provide a streamlined workflow for designing and executing RL missions. Key features include: (1) high-level methods for retrieving orbital state information, managing agents, and configuring scenarios; (2) flexibility to define single-agent or multi-agent missions, supporting tasks such as station-keeping, orbital transfers, and collision avoidance; and (3) compatibility with PettingZoo, enabling seamless integration with existing RL workflows.
4.3 Addressing Key Challenges
OrbitZoo directly addresses the challenges outlined in Section 3:
Bridging the Reality Gap. OrbitZoo leverages Orekit’s high-fidelity propagator to model perturbative forces like atmospheric drag, solar radiation pressure, and third-body effects. A key feature is its flexibility in replicating real-world data, such as Starlink satellite orbits. This is enabled by its modular architecture, allowing precise customization of orbital parameters, perturbations, and thrust models. For instance, by optimizing parameters such as drag and reflection coefficients using Bayesian techniques, OrbitZoo can replicate Starlink’s orbital trajectories as provided by ephemeris data. This feature bridges the reality gap by ensuring that the simulated environment closely resembles the dynamics of real-world systems. As demonstrated in Section 5, the propagated trajectories generated by OrbitZoo achieve a low Root Mean Square Error (RMSE) when compared to Starlink ephemeris data, validating its suitability for high-fidelity reinforcement learning applications.
Realistic Thrust Modeling. The inclusion of polar thrust parametrization and configurable action spaces allows agents to perform realistic and efficient maneuvers, overcoming limitations in traditional thrust models.
Multi-Agent Coordination. OrbitZoo supports decentralized and centralized multi-agent training, enabling the simulation of satellite constellations and cooperative missions. Generalized advantage estimation (GAE) and PPO algorithms ensure stability and scalability in multi-agent setups.
Interactive Visualization for Debugging and Analysis. The visualization capabilities of OrbitZoo allow users to inspect agent behavior and system dynamics in real-time, making it essential for debugging and validating agent policies in a complex, partially observable environment.
5 Experiments and Results
To evaluate the effectiveness of OrbitZoo in modeling diverse orbital missions and supporting reinforcement learning (RL) methods, we conducted a series of experiments. These include a single-agent Hohmann transfer maneuver, a multi-agent geostationary orbit (GEO) constellation coordination problem, and a validation experiment comparing OrbitZoo’s simulations against real-world Starlink ephemeris data. These scenarios were chosen to demonstrate OrbitZoo’s ability to bridge the reality gap, enable realistic control tasks, and support scalable multi-agent coordination.
The main goal of these experiments was to validate OrbitZoo’s ability to model high-fidelity orbital dynamics, demonstrate the generalization of learned policies, and provide a robust environment for RL approaches, rather than aiming for state-of-the-art performance in each task. The following sections outline the setup and results of each experiment.
5.1 Single-Agent Hohmann Maneuver
The Hohmann transfer experiment was chosen as a benchmark to evaluate OrbitZoo’s ability to support reinforcement learning (RL) in a high-fidelity orbital dynamics environment. As a classic problem in orbital mechanics, the Hohmann transfer is analytically solvable and provides a clear reference for comparing RL-derived solutions with theoretical optima [22].
Setup
The agent observes the satellite’s current orbital state through equinoctial coordinates and receives a reward based on the reduction of transfer error while minimizing fuel consumption. The action space corresponds to polar thrust parameters $(T,\theta,\phi)$, where $T$ is the thrust magnitude and $(\theta,\phi)$ define the thrust direction. The environment incorporates perturbative forces, including drag.
Results
The RL agent learned near-optimal strategies for the Hohmann transfer, closely matching theoretical values for the semi-major axis. However, deviations in other orbital elements, such as inclination, occurred due to thrust misalignments. When tested with novel perturbative forces, the agent adapted and reached the target. These results validate OrbitZoo’s ability to model complex orbital dynamics and highlight opportunities to refine RL policy design and reward functions. Figure 12 shows the optimized transfer trajectory, achieving the target orbit with minimal fuel use. The trajectory closely matches the theoretical solution (Figure 3), demonstrating OrbitZoo’s high-fidelity modeling. Details of the experiment setup, training hyperparameters, and further quantitative results can be found in Appendix 0.D.2.
Figure 3: The L2 distance error between the optimal maneuver and Experiment 2 on the target orbit remains small over a large orbital period, with the larger error resulting from small inclination changes. However, the agents were trained to minimize differences in equinoctial elements rather than Euclidean distance.
5.2 Validation Against Real-World Data
To evaluate OrbitZoo’s ability to bridge the reality gap, we configured the environment to simulate Starlink satellites and compared its output to publicly available ephemeris data in Space-Track.
Setup
Four Starlink satellites with publicly available data were selected for the validation experiment. Bayesian optimization was employed to tune parameters such as the drag coefficient, reflection coefficient, and satellite radius to match the ephemeris data. The RMSE between the OrbitZoo-propagated trajectories and the ephemeris data was used as the primary metric for evaluation.
Results
As summarized in Table 2, OrbitZoo achieved varying levels of accuracy across 31 satellites, with mean RMSE values ranging from 24.14 meters to 1924.90 meters over 16.6-hour propagation. While some satellites closely matched the Starlink ephemeris data, others showed significant deviations, likely due to limited information on their physical properties. Figure 4 illustrates the residuals between the propagated and observed trajectories, demonstrating the accuracy of OrbitZoo’s physical modeling for the relevant horizon of two hours.
Table 2: Mean RMSE values for three groups of satellites (a total of 31 satellites), separated based on an RMSE threshold. The "Low RMSE" group includes satellites with RMSE values below 50, the "Medium RMSE" group includes satellites with RMSE values between 50 and 100, and the "High RMSE" group includes satellites with RMSE values above 100. These RMSE values are derived from a 16.6 hour propagation (1000 steps).
Group Mean RMSE (meters)
Low RMSE 24.14
Medium RMSE 83.75
High RMSE 1924.90
Figure 4: Residuals between OrbitZoo-propagated trajectories and Starlink ephemeris data for satellite 44748. The residuals remain low over the validation interval, confirming OrbitZoo’s high fidelity.
5.3 Multi-Agent GEO Constellation Coordination
This experiment demonstrates OrbitZoo’s ability to simulate and train multi-agent systems for satellite constellation management in geostationary orbit (GEO).
Setup
A four-satellite constellation in GEO aimed to maintain equal angular separation and altitude while minimizing fuel use through small thrusts. Each agent observes the anomalies of other satellites, its own semi-major axis and eccentricity, and outputs thrust commands via a decentralized policy. Training used PPO with generalized advantage estimation (GAE).
Results
Figure 5: Visualization of the initial random constellation and its evolution over nearly 4 days, with the purple circle representing the GEO orbit. After this period, the agents exhibit more distinct anomalies while maintaining proximity to the nominal orbit.
Figure 5 shows a view of the OrbitZoo interface of the constellation’s configuration after 4 days. The agents successfully maintained angular separation while minimizing fuel consumption. The policies are also able generalize to unseen perturbations, such as third body forces (Sun and Moon), solar radiation pressure and drag. Further results and a detailed explanation can be found in Appendix 0.D.5.
5.4 Additional Experiments
For completeness, we conducted a chase target experiment, where a satellite was tasked with pursuing a moving target in a higher orbit, and a single-agent collision avoidance experiment, where a satellite had to evade debris in LEO. The setup and results of these experiments are detailed in Appendix 0.D.3 and Appendix 0.D.4, respectively.
Conclusions
We introduced OrbitZoo, a high-fidelity reinforcement learning (RL) environment for orbital dynamics that integrates with an industry-standard physics engine (Orekit). Unlike prior RL environments with simplified models, OrbitZoo incorporates realistic perturbative forces, flexible coordinate representations, and multi-agent support. It enables data-driven decision-making for space missions, bridging the reality gap between simulation and real-world orbital operations. Through a series of experiments—including a Hohmann transfer, a GEO constellation coordination task, and validation against real-world Starlink data—we demonstrated that RL policies trained in OrbitZoo can achieve near-optimal control strategies while aligning with physical satellite behavior. These results establish OrbitZoo as a credible benchmark for RL-based autonomy in space operations. Beyond RL, OrbitZoo serves as an open-source platform for researchers in aerospace engineering, space operations, and machine learning. Future work includes expanding real-world validation with more satellite datasets, integrating real-time onboard learning, and enhancing adaptive multi-agent coordination. By enabling realistic and reproducible RL research for space applications, OrbitZoo lays the groundwork for trustworthy AI-driven autonomy in future space missions.
Impact Statement
Reinforcement learning (RL) has shown promise in complex decision-making but struggles with real-world orbital dynamics due to the reality gap. OrbitZoo bridges this gap by providing a high-fidelity, modular RL simulation framework integrated with Orekit, supporting both single-agent and multi-agent tasks.
This work advances RL research in high-dimensional continuous control for space operations by incorporating perturbative forces, uncertainties, and ephemeris-based validation. Our experiments show that RL policies trained in OrbitZoo closely match real-world Starlink trajectories, establishing a benchmark for autonomous space systems.
Appendix 0.A Orbital Mechanics
This section provides an overview of key orbital mechanics concepts to support a clearer understanding of the concepts presented throughout the paper. We begin by explaining the most commonly used coordinate systems, then discuss how body propagation is usually managed, and conclude with our approach to realistic thrust representation.
0.A.1 Coordinate Systems
Orbital dynamics is typically described using different coordinate systems, each offering unique advantages depending on the problem at hand. Considering a large central body (such as Earth) as an inertial frame, Cartesian coordinates provide a straightforward and intuitive representation of a body’s position and velocity in space as $r=(x,y,z)$ and $\dot{r}=(\dot{x},\dot{y},\dot{z})$, respectively, as shown in Figure 6. While useful for direct numerical computations, such as altitude or distance between bodies, this state rapidly changes and often lacks the deeper insights into the actual orbital shape and orientation, which many missions focus on, such as the closest point (periapsis) or farthest point (apoapsis) from the orbiting body.
Figure 6: Representation of Cartesian position ($r$) and velocity ($\dot{r}$).
Keplerian elements, as Figure 7 shows, describe the orbital motion using six parameters: the semi-major axis $(a)$, eccentricity $(e)$, inclination $(i)$, argument of perigee $(\omega)$, longitude of the ascending node $(\Omega)$, and anomaly (which can be represented in three ways, as seen in Figure 8). However, the mean anomaly is commonly used due to its linear evolution over time.
Figure 7: Representation of Keplerian Orbital Elements. Here, it is assumed that the orbit is perfectly circular ($e=0$), and consequently, the semi-major axis corresponds to the Euclidean distance to the primary focus ($M$).
Figure 8: Types of anomaly. A body $S$ orbits a central body $C$ in an elliptical orbit with semi-major axis $a$. The true anomaly ($\theta$) is the actual angular position of $S$ measured from the periapsis (closest point to $C$). The mean anomaly ($M$) is the angle assuming uniform motion over the same orbital period (represented by body $S^{\prime}$). Eccentric anomaly ($E$) is an intermediate angle used to relate $M$ and $\theta$, representing the position of $S$ on the circular orbit ($Q$).
These elements directly relate to the orbital geometry and are particularly useful for characterizing two-body motion. However, they can become undefined or poorly behaved in special cases, such as circular or equatorial orbits. These scenarios, referred to as singularities, pose significant challenges for numerical integration and optimization problems. Equinoctial elements are therefore preferred for RL, as they avoid these singularities by using five parameters (plus anomaly): $(a,e_{x},e_{y},h_{x},h_{y})$, where $e_{x}$ and $e_{y}$ represent components of the eccentricity vector, and $h_{x}$ and $h_{y}$ describe the inclination vector.
0.A.2 Thrust Representation
Spacecraft control often employs local reference frames (centered on the body) for thrust actions due to their intuitive representation, such as RSW. Given a satellite’s Cartesian position $r$ and velocity $\dot{r}$, it is straightforward to compute the radial ($\hat{r}$), cross-track ($\hat{w}$) and along-track ($\hat{s}$) unit vectors:
$\hat{r}=\frac{r}{\|r\|}\quad\quad\hat{w}=\frac{r\times\dot{r}}{\|r\times\dot{r}\|}\quad\quad\hat{s}=\hat{w}\times\hat{r},$ (1)
where $\times$ represents the cross product.
In this approach, the thrust usually consists of a vector $\mathbf{T}_{\text{RSW}}=(T_{R},T_{S},T_{W})$ that represents the thrust magnitude on each axis of the RSW frame. The action space is then limited to a maximum thrust magnitude $T_{\text{max}}$ on each component: $\mathbf{T}_{\text{RSW}}\in[-T_{\text{max}},T_{\text{max}}]^{3}$.
A more realistic and versatile approach to modeling the thrust action space is to adopt a polar parametrization, representing the thrust as $\mathbf{T}=(T,\theta,\phi)$. A visual comparison can be seen in Figure 9. This parameterization limits the thrust to the action space $\mathbf{T}\leq(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}})$. When $\theta_{\text{max}}=\pi$ rad and $\phi_{\text{max}}=2\pi$ rad, the satellite can apply thrust in any possible direction. This parameterization naturally constrains the thrust to a cone-shaped region, providing a physically realistic limit on the directions in which the satellite can apply force, offering a more controlled and flexible representation of the satellite’s maneuvering capabilities. Finally, the thrust vector in the RSW frame can be obtained via the following transformation:
$\mathbf{T}_{\text{RSW}}=T(\cos{\theta}\hat{s}+\sin{\theta}(\cos{\phi}\hat{r}+\sin{\phi}\hat{w})).$ (2)
Now consider that the spacecraft has a thrust system that produces a force $\mathbf{T}\in\mathbb{R}^{3}$. Similarly to the gravitational acceleration, the actual acceleration provoked by this thrust system comes from Newton’s second law of motion:
$\mathbf{T}=ma_{\text{thrust}}\Leftrightarrow a_{\text{thrust}}=\frac{\mathbf{T}}{m}.$ (3)
This expression indicates that propulsive capability is inversely related to the spacecraft mass; hence, as propellant is expended, the spacecraft is capable of higher acceleration values when using the same force. The rate at which mass is lost is given by:
$\dot{m}=-\frac{\|\mathbf{T}\|}{I_{\text{sp}}g},$ (4)
where $I_{\text{sp}}$ is the specific impulse (a measure of the efficiency of the propulsion system) and $g$ is standard gravity. A higher $I_{\text{sp}}$ value means that the spacecraft can generate more thrust per unit of propellant mass, which makes them more efficient by saving fuel.
(a) Cartesian parametrization.
(b) Polar parametrization.
Figure 9: Comparison of Cartesian and Polar parameterizations of the thrust vector $\mathbf{T}$.
0.A.3 Propagation
The motion of celestial bodies is governed by Newton’s law of universal gravitation, which states that each body exerts a gravitational force on every other body. For a two-body system, the force acting on a satellite or debris with mass $m$ due to a central body with mass $M$ is given by:
$F=-\frac{GMm}{\|r\|^{2}}\hat{r},$ (5)
[…]
$\ddot{r}=-\frac{GM}{\|r\|^{2}}\hat{r}=-\frac{\mu_{E}}{\|r\|^{2}}\hat{r},$ (6)
where $\mu_{E}$ is called the standard gravitational parameter of Earth.
Although this model forms the foundation of orbital mechanics and is commonly used in RL, analytical models like Simplified General Perturbations (SGP) incorporate additional perturbative forces – such as atmospheric drag, solar radiation pressure, and Earth’s oblateness – to provide a fast and relatively accurate method of propagating bodies.
Numerical propagators provide a powerful alternative to SGP by solving the equations of motion through numerical integration, enabling precise modeling of complex orbital dynamics. Unlike SGP, numerical propagators can incorporate a wide range of perturbative forces, including higher-order gravitational harmonics, third-body effects, atmospheric drag, and solar radiation pressure. This flexibility allows them to handle scenarios requiring high precision, such as long-term orbit predictions and mission-critical maneuvers.
Assuming the state of a spacecraft is characterized by its Cartesian position $r$, velocity $\dot{r}$ and mass $m$, the state vector $s=(r,\dot{r},m)\in\mathbb{R}^{7}$ can be propagated in time using integration methods. The Runge-Kutta (RK4) method is a numerical integration technique used to solve first-order differential equations, which can be used to approximate the unknown function $s$ dependent on time $t$. Since the state corresponds to a second-order system, it can be rewritten as a first-order system:
$f(t,s)=\frac{d}{dt}s=\frac{d}{dt}\begin{bmatrix}r\\ \dot{r}\\ m\end{bmatrix}=\begin{bmatrix}\dot{r}\\ \ddot{r}\\ \dot{m}\end{bmatrix}=\begin{bmatrix}\dot{r}\\ -\frac{\mu_{E}}{\|r\|^{2}}\hat{r}+\frac{T}{m}+a_{\text{env}}\\ -\frac{\|T\|}{I_{\text{sp}}g}\end{bmatrix}.$ (7)
By knowing the state of the spacecraft at a given moment, we create the initial conditions $t_{0}$ and $s_{0}$. To get an approximation of the state after a step size $\Delta t$, we define:
$s_{\Delta t}=s_{0}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4}),$ (8)
where
$k_{1}$ $=f(t_{0},s_{0}),$
$k_{2}$ $=f\left(t_{0}+\frac{\Delta t}{2},s_{0}+\Delta t\frac{k_{1}}{2}\right),$
$k_{3}$ $=f\left(t_{0}+\frac{\Delta t}{2},s_{0}+\Delta t\frac{k_{2}}{2}\right),$
$k_{4}$ $=f\left(t_{0}+\Delta t,s_{0}+\Delta tk_{3}\right).$
Appendix 0.B Mathematical Definitions
For completeness, we provide the mathematical details of the methods referred to in the main paper.
0.B.1 Proximal Policy Optimization (PPO)
The actor loss in PPO is defined based on the ratio $r_{t}(\theta)$, which measures the probability of taking an action $a_{t}$ under the current policy compared to the old policy:
$r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}.$ (9)
The clipped objective ensures stability during training by preventing excessively large updates to the policy:
$\min_{\theta}\mathbb{E}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right],$ (10)
where $\epsilon$ controls the trust region, and $\hat{A}_{t}$ is the advantage function. The advantage $\hat{A}_{t}$ is computed recursively using the temporal difference (TD) error:
$\delta_{t}=r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t}),$ (11)
$\hat{A}_{t}=\delta_{t}+\gamma\lambda\cdot\hat{A}_{t+1},$ (12)
where $\lambda$ is a decay factor and $V_{\phi}(s_{t})$ is the value function approximated by the critic network.
0.B.2 Generalized Advantage Estimation (GAE)
Generalized Advantage Estimation (GAE) provides a flexible way to compute the advantage by balancing bias and variance through the $\lambda$ parameter:
$\hat{A}_{t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}.$ (13)
For long episodes, such as those encountered in orbital dynamics, GAE effectively stabilizes training by reducing variance in advantage estimation.
Appendix 0.C Computational Performance
Real orbital systems can consist of hundreds or even thousands of bodies, each requiring propagation at every step. Additionally, different forces may act on different bodies, and each body can have unique properties such as shape, attitude, or drag/reflection coefficients.
OrbitZoo imposes no strict limit on the number of bodies in a system, with scalability constrained only by the available hardware (0.C.1). However, it is essential to assess how increasing the number of bodies and introducing complex dynamics impact simulation speed and parallelization efficiency. These aspects are examined in the following subsections (0.C.2 and 0.C.3).
Although not yet implemented, OrbitZoo experiments could benefit from libraries mentioned in 3.4 to accelerate training.
0.C.1 Hardware Specifications
OrbitZoo supports systems of varying sizes and complexity, making hardware requirements dependent on the specific system being modeled, particularly in terms of CPU power and memory. For the following experiments and evaluations, the hardware used is detailed in Table 3, with the GPU utilized solely for training RL agents.
Table 3: Hardware specifications.
Hardware Specification
CPU Intel(R) Core(TM) i3-8100 CPU @3.60 Hz, 3600 Mhz, 4 Cores, 4 Logical Processors
GPU NVIDIA GeForce GTX 1050 Ti
RAM 16.0 GB, 2933 Mhz
0.C.2 Scalability
One of the simplest metrics for assessing scalability is simulation speed. Specifically, the time required to perform a single propagation step for all bodies in the system. Since multiple factors influence this speed, we focus on addressing the following questions:
  • How does the simulation speed evolve with the addition of bodies?
  • How does the simulation speed evolve with the addition of forces acting on the bodies (more realism)?
To address the first question, we evaluate system performance when propagating 1, 5, 10, 100, 1000, and 10000 bodies. For the second question, we analyze simulation speed starting with a simple Newtonian gravity model assuming a perfectly spherical central body. We then introduce a more complex gravity model using spherical harmonics (HolmesFeatherstone) and further incorporate perturbative forces, including third-body effects (Sun and Moon), solar radiation pressure, and drag.
Results in Figure 10 show that OrbitZoo exhibits approximately $O(n)$ time complexity per step, where $n$ is the number of bodies in the system. This trend becomes more evident as $n$ increases. However, for smaller systems (1 to 10 bodies), the effects of parallelization are noticeable, as the hardware and OrbitZoo distribute computations across multiple threads. Additionally, the inclusion of perturbative forces increases the step time in a roughly linear manner.
Scalability and performance also depend on user-specific implementations. For example, in large LEO satellite constellations, drag is a crucial perturbative force for realistic trajectory modeling. However, drag computation relies on body shape, which is assumed to be unique for each satellite by default. As a result, each propagator stores its own atmospheric data, significantly increasing memory usage. By assuming uniform body characteristics (e.g., identical shapes), a shared force instance can be used across propagators, drastically reducing memory consumption.
Figure 10: Time per step (in seconds) by number of bodies and active forces. Each point represents an average of 100 propagation steps of 10 seconds.
0.C.3 Parallelization
As shown in Figure 1, OrbitZoo integrates with two key external components: PettingZoo and Orekit. PettingZoo provides a framework for sequential or parallelized step computations, while Orekit handles orbital dynamics within each step. Since orbital MARL missions assume all agents act simultaneously rather than in turns, OrbitZoo propagates systems in a parallel manner.
OrbitZoo leverages PettingZoo’s parallelization by implementing the ParallelEnv class. To assess the impact of this choice, we compared simulation times between parallel and sequential step computations. Specifically, we measured how much longer sequential execution takes relative to parallel execution as the number of bodies increases (1, 5, 10, 100, 1000, and 10000 bodies).
Figure 11 shows that using a sequential propagation architecture instead of a parallel one leads to a linear increase in step computation time with the number of bodies. For instance, with 100 bodies, sequential propagation takes nearly 0.1 seconds longer per step than parallel propagation, while with 10000 bodies, the difference exceeds 1 second per step. Given that a single RL episode can consist of hundreds or thousands of steps, this improvement significantly enhances simulation efficiency.
Figure 11: Additional time per step (in seconds) when using sequential computation instead of parallel, by the number of bodies. Each point represents an average of the difference of 100 propagation steps of 10 seconds, between the sequential and parallel step times.
Orekit also offers tools for parallelizing propagations, such as the PropagatorsParallelizer class, which creates new threads to speed up computations. OrbitZoo could benefit from these tools to further improve simulation time. However, this has not been implemented yet, as parallelization in Java (native to Orekit) is handled differently in the Python wrapper. This discrepancy presents a challenge, but it will be an area of future research. As of now, the individual propagation of bodies remains the primary bottleneck for parallelization and scalability within OrbitZoo.
Appendix 0.D Experiments
To evaluate OrbitZoo’s capabilities, we designed a series of missions utilizing a state-of-the-art algorithm (0.D.1). The first two missions (0.D.2, 0.D.3) focus on orbital transfers (OTs), where agents learn strategies to reach static and dynamic targets, respectively, while accounting for real-world factors such as mass loss due to fuel consumption. The third mission (0.D.4) involves a satellite learning to minimize collision probability with debris in the days leading up to the predicted closest approach, all while maintaining its initial orbit. Lastly, the fourth mission (0.D.5) is a multi-agent scenario in which four satellites learn to distribute themselves evenly along a GEO orbit.
These missions capture key aspects of both orbital dynamics and reinforcement learning (RL). Specifically, in terms of realism, we model various bodies, including drifters and satellites, with distinct characteristics such as radius, dry and fuel mass, and propulsion systems featuring different specific impulses and thrust magnitudes. From an orbital dynamics perspective, we train agents using simplified dynamics and evaluate their learned policies in more complex environments to assess generalization, with real perturbations present. Additionally, we leverage multiple coordinate systems and OrbitZoo’s built-in methods to facilitate efficient implementations across all missions.
For each experiment, we first provide a brief definition, highlighting key aspects of the mission. We then describe the environment setup, detailing the observation and action spaces, body characteristics, and reward functions. Finally, we present the results, assessing the generalization capabilities of each policy.
0.D.1 Learning Algorithm
The chosen algorithm for all missions was PPO, since it is a state-of-the-art RL algorithm in environments with continuous action spaces, such as in the experiments carried out in this paper. The implementation was inspired by [8], which offers a minimal PPO setup for discrete and continuous action spaces. However, several modifications were introduced to enhance learning performance, as recommended in [3]. These include the use of generalized advantage estimation (GAE) instead of Monte Carlo estimation for calculating advantages, recalculating advantages at each epoch, and employing batch-based training. Other improvements include changes in the network architecture (with input normalization, tanh activation functions, and smaller initial weights), together with hyperparameter tuning and performance adjustments (such as calculation of expected returns only at training time).
The overall structure of the actor and critic networks used throughout the experiments is similar, with two hidden layers and Tanh activation functions, as represented in Table 4. No extensive research was made to find optimal hyperparameters.
Table 4: Network Structure of PPO Actor and Critic.
Layer Actor Network (Input: state_dim_actor) Critic Network (Input: state_dim_critic)
Input BatchNorm1d(state_dim_actor) BatchNorm1d(state_dim_critic)
Hidden Layer 1 Linear(state_dim_actor, 500), Tanh Linear(state_dim_critic, 500), Tanh
Hidden Layer 2 Linear(500, 450), Tanh Linear(500, 450), Tanh
Output Linear(450, action_dim), Tanh Linear(450, 1)
0.D.2 Hohmann Maneuver
The Hohmann transfer experiment [22] was chosen due to its well-established relevance in orbital mechanics and its suitability as a benchmark for evaluating reinforcement learning frameworks in this domain. As one of the most fuel-efficient orbital transfer maneuvers, it provides a clear and analytically solvable problem that allows for direct comparison between RL-derived solutions and theoretical optima. Moreover, the Hohmann transfer incorporates key aspects of orbital dynamics, such as thrust application, trajectory optimization, and state transitions, making it an ideal testbed to validate the realism and effectiveness of OrbitZoo’s high-fidelity environment.
0.D.2.1 Definition
Consider a spacecraft on a nearly circular orbit, where its semi-major axis approximately corresponds to the distance of the spacecraft to the primary focus ($R$). If this spacecraft has the objective of ending up on an orbit with an increased distance of $R^{\prime}$ while maintaining all other elements constant, the Hohmann transfer maneuver can achieve this in a highly fuel-efficient manner by applying two very specific impulsive thrusts in the along-track direction ($\Delta V$ and $\Delta V^{\prime}$[51]. The first establishes the transfer orbit (with a semi-major axis $R<a_{H}<R^{\prime}$), and the second adjusts the elements to match the target orbit. When considering instantaneous impulses and conservation of energy, these changes in velocity can be easily calculated:
$\Delta V=\sqrt{\frac{\mu}{R}}\left(\sqrt{\frac{2R^{\prime}}{R+R^{\prime}}}-1\right)\quad\quad\Delta V^{\prime}=\sqrt{\frac{\mu}{R^{\prime}}}\left(1-\sqrt{\frac{2R}{R+R^{\prime}}}\right),$ (14)
where $\mu$ is the standard gravitational parameter of Earth ($\mu=GM\approx 3.986\text{ m}^{3}\text{s}^{-2}$), and $R$ and $R^{\prime}$ are the radii of the departure and arrival orbit, respectively. According to the Tsiolkovsky rocket equation [50], a body with total mass $m_{0}$ and $m_{0}-m_{f}$ of fuel mass, and with thrusting capabilities that have a specific impulse $I_{sp}$, can only perform the Hohmann maneuver if $\Delta V+\Delta V^{\prime}\leq I_{sp}g_{0}\ln{\frac{m_{0}}{m_{f}}}$, where $g_{0}$ is standard gravity. In a real scenario, the force ($F_{1}$) that is required to change the velocity by $\Delta V$ depends on the mass of the body and the time that we are applying that force ($\Delta t$):
$F_{1}=\frac{\Delta V\times m_{0}}{\Delta t}.$ (15)
After this maneuver, the body loses some mass ($\dot{m}$) that is inversely proportional to the specific impulse of the thrust:
$\dot{m}=\frac{F_{1}}{I_{sp}g_{0}}.$ (16)
Although small, this difference in mass will impact the force ($F_{2}$) that should be applied to change the velocity by $\Delta V^{\prime}$:
$F_{2}=\frac{\Delta V^{\prime}\times(m_{0}-\dot{m})}{\Delta t}.$ (17)
Finally, the time between $\Delta V$ and $\Delta V^{\prime}$ (transfer time, $t_{H}$) is given by Kepler’s third law:
$t_{H}=\frac{1}{2}\sqrt{\frac{4\pi^{2}a_{H}^{3}}{\mu}},$ (18)
where usually it is assumed that $a_{H}=(R+R^{\prime})/2$. A representation is shown in Figure 12.
Figure 12: Representation of the Hohmann Maneuver. A spacecraft is initially in an orbit with semi-major axis $R$. It applies a thrust $\Delta V$ which positions it on an eccentric orbit. When the spacecraft reaches apoapsis, it performs a second thrust $\Delta V^{\prime}$ which positions it on the intended red orbit, with a semi-major axis of $R^{\prime}$.
0.D.2.2 Environment Setup
We define a spacecraft with a mass of $m_{f}=200$ kg and $50$ kg of fuel, starting on a near-circular orbit with equinoctial elements $(a,e_{x},e_{y},h_{x},h_{y})=(2000+R_{E}\text{ km},0.007,0.006,0.041,0.015)$ and ending up on $(a,e_{x},e_{y},h_{x},h_{y})=(2030+R_{E}\text{ km},0.007,0.006,0.041,0.015)$, where $R_{E}=6378$ is the radius of Earth. Therefore, the objective is to raise the semi-major axis by 30 km.
According to these values, using the Hohmann transfer, the optimal maneuver requires $\Delta V=\Delta V^{\prime}=6.16$ m/s and a transfer time of $t_{H}=3826.1$ s. Therefore, we define that each episode contains 1000 steps of 5 seconds, giving extra time for the agents to reach the target. Instead of assuming instantaneous burns, which are unrealistic, the thrust is applied for the entire step (5 seconds). For this interval, the optimal forces are $F_{1}=308$ N and $F_{2}=307.9$ N in the along-track direction. The spacecraft is equipped with a thruster that provides a specific impulse of $I_{sp}=310$ s and is capable of performing thrusts in every direction.
For simplicity, we consider a Newtonian attraction model with no perturbations, and the agents to always start in the same anomaly with little uncertainty. Tthe agent completes its objective if it reaches the target orbit within a tolerance for every equinoctial element: (±$(\pm 100\text{ m},\pm 0.005,\pm 0.005,\pm 0.001,\pm 0.001)$.
The observation $o_{t}$ is defined as a vector comprising the current equinoctial elements $s_{t}=(a_{t},e_{xt},e_{yt},h_{xt},h_{yt})$, the mean anomaly $M_{t}$, and the remaining fuel $f_{t}$: $o_{t}=(s_{t},M_{t},f_{t})\in\mathbb{R}^{7}$.
To account for the requirement that thrust should be applied only at specific moments during the episode rather than continuously, an additional component is included in the action space: $a_{t}=(T,\theta,\phi,\delta)$. Here, $\delta$ represents the decision to apply thrust, where values close to 1 or 0 indicate a high certainty to apply or retain thrust, respectively. The action space is limited by $(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}},\delta_{\text{max}})=(500,\pi,2\pi,1)$.
The reward function is designed to encourage the agent to apply large but accurate thrusts. First, the absolute difference between the current orbit at time $t$ and the target orbit, $\hat{s}$, is defined as $\Delta s_{t}=|s_{t}-\hat{s}|$. Next, the scalar progress is computed as $P_{t}=w^{T}\cdot(\Delta s_{t-1}-\Delta s_{t})/s_{t}$, where $w$ is a vector of constant weights assigned to each orbital element. The reward function is then formulated as:
$r_{t}=\mathbbm{1}_{\delta>0.5}\left(\alpha_{1}\frac{T}{T_{\text{max}}}P_{t}-\alpha_{2}\frac{\theta}{\theta_{\text{max}}}\right),$ (19)
where $\mathbbm{1}_{\delta>0.5}$ is an indicator function that activates the reward only when the thrust decision $\delta$ exceeds 0.5, and $\alpha_{1}$ and $\alpha_{2}$ are the importance given to the progress, and actions in the along-track direction, respectively.
Table 5: Hohmann maneuver: Training hyperparameters.
Parameter Value
Actor learning rate 0.0001
Critic learning rate 0.001
GAE $\lambda$ 0.95
Epochs 5
Discount factor ($\gamma$) 0.99
Clip ($\epsilon$) 0.1
Initial Standard Deviation 0.4
Standard Deviation Decay Rate 40000 steps
Standard Deviation Decay Amount 0.05
Minimum Standard Deviation 0.05
Experiences for Training 4096
Batch Size 64
0.D.2.3 Results
The agent training hyperparameters are shown in Table 5. Additionally, two experiments were conducted to analyze the agent’s behavior, with two different reward functions as shown in Table 6. However, there was no extensive exploration to find optimal coefficients. For the second experiment, the generalization of the agent was tested by switching from a Newtonian attraction model to a spherical harmonic gravity, and using third body forces (Sun and Moon), solar radiation pressure and drag force.
Table 6: Values of $\alpha_{1}$ and $\alpha_{2}$ used in each experiment to analyze the agent’s behavior. The first experiment does not penalize thrust deviations from the along-track direction, while the second experiment includes a penalty for such deviations.
Experiment $\boldsymbol{\alpha_{1}}$ $\boldsymbol{\alpha_{2}}$
Experiment 1 $1$ 00
Experiment 2 $1$ $0.5$
The agent was trained until it successfully adopted a strategy to achieve the target orbit within the specified tolerance. Figure 13 illustrates that in both experiments, the agent achieved the target semi-major axis while also correcting some other orbital elements. However, in Experiment 1, the agent failed to bring the $h_{y}$ component within the tolerance. In contrast, the results of the second experiment show a closer approximation to the optimal maneuver. Notably, when using more realistic dynamics by adding perturbative forces, the agent is also able to reach the target orbit, even with a constant change in orbital parameters created by these perturbations.
Nevertheless, both experiments demonstrate that the agent incorrectly learned to apply thrusts that alter the inclination, as shown in Figure 14, with thrusts being applied in the radial ($r$) and cross-track ($w$) directions. Finally, Figure 15 illustrates that the agent’s decision policy $\delta$ is nearly optimal, with thrust activations occurring both at the beginning of the episode and near $t_{H}$.
Overall, this experiment demonstrates that even when trained without perturbations, the policy successfully generalizes to real dynamics, adapting to complex gravity fields, third-body forces, solar radiation pressure, and atmospheric drag. This is particularly significant in LEO, where gravity and drag are the dominant influences on a satellite’s trajectory.
Figure 13: Evolution of equinoctial elements ($\Delta a,e_{x},e_{y},h_{x},h_{y}$) in the Hohmann maneuver for the first experiment (red), second experiment (yellow), second experiment with realistic perturbations (purple) and optimal maneuver (green). $\Delta a$ corresponds to the absolute difference between the target and current semi-major axis.
Figure 14: Comparison of the applied thrust ($r,s,w$), in Newtons, for the first experiment (red), second experiment (yellow), second experiment with realistic perturbations (purple) and optimal maneuver (green).
Figure 15: Comparison of the agents decisions ($\delta$) to apply thrust during a full episode.
0.D.3 Chase Target
0.D.3.1 Definition
Orbital transfer missions typically involve placing a satellite into a specific stationary orbit, as seen in the Hohmann transfer, or performing station-keeping maneuvers to maintain a nominal orbit. However, targets in such missions can also be nonstationary, as in scenarios where the objective is to rendezvous with moving bodies such as debris or active satellites. Many missions focus on approaching dynamic targets, like comets, particularly when the goal is scientific observation and study.
0.D.3.2 Environment Setup
Table 7: Chase Target: Training hyperparameters.
Parameter Value
Actor learning rate 0.00001
Critic learning rate 0.0001
GAE $\lambda$ 0.95
Epochs 5
Discount factor ($\gamma$) 0.99
Clip ($\epsilon$) 0.1
Initial Standard Deviation 0.4
Standard Deviation Decay Rate 10000 steps
Standard Deviation Decay Amount 0.05
Minimum Standard Deviation 0.05
Experiences for Training 4096
Batch Size 64
In this experiment, we have an agent/satellite, called follower, that starts on an orbit characterized by the following Keplerian elements $(a_{F},e_{F},i_{F},\omega_{F},\Omega_{F},M_{F})=(10000+R_{E}\text{ km},0.1.5.0^{\circ},10.0^{\circ}.10.0^{\circ}.10.0^{\circ})$. Additionally, there is a non-maneuverable body, called leader, that starts on an orbit with a much larger altitude and less inclination than the follower: $(a_{L},e_{L},i_{L},\omega_{L},\Omega_{L},M_{L})=(40000+R_{E}\text{ km},0.001.5.0^{\circ},10.0^{\circ}.10.0^{\circ}.10.0^{\circ})$. To avoid singularities, we use equinoctial parameters to represent the orbits of the follower $(a_{F},e_{x_{F}},e_{y_{F}},h_{x_{F}},h_{y_{F}})$ and the leader $(a_{L},e_{x_{L}},e_{y_{L}},h_{x_{L}},h_{y_{L}})$. The objective is for the follower to approximate the leader.
The satellite is characterized by being spherical with a radius of 5 m, a mass of 650 kg (from which 150 kg are fuel), and has a propulsion system containing a specific impulse of $I_{sp}=3000$ s. Finally, the dynamics for training are composed of Newtonian attraction without additional perturbations, and each episode contains steps of 70 seconds.
Since the orbital elements of the leader remain static throughout the episode (besides the anomaly $M_{L}$), the observation space of the agent becomes smaller, as it does not need to know those elements: $o_{t}=(a_{F},e_{x_{F}},e_{y_{F}},h_{x_{F}},h_{y_{F}},M_{F},M_{L},f_{t})$, where $f_{t}$ represents the current fuel.
The satellite thrust system is capable of performing thrusts in any direction with a maximum magnitude of 30 N, that is, $(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}})=(30,\pi,2\pi)$.
The reward function aims to minimize the error between each equinoctial element, including the anomaly. We first define the difference between each element $e$ as: $\Delta e=e_{L}-e_{F},\forall{e}\in\{a,e_{x},e_{y},h_{x},h_{y},M\}$. The final reward function weights each absolute difference by a factor $\alpha_{e}$:
$r_{t}=-\left(\frac{|\Delta a|}{a_{L}},|\Delta e_{x}|,\Delta e_{y}|,|\Delta h_{x}|,|\Delta h_{y}|,|\text{atan2}(\sin\Delta M,\cos\Delta M)|\right)\cdot\alpha^{T},$ (20)
where $\alpha=(\alpha_{a},\alpha_{e_{x}},\alpha_{e_{y}},\alpha_{h_{x}},\alpha_{h_{y}},\alpha_{M})$. The error for the semi-major axis is normalized by the target element (leader) due of its large scale, and the error in the anomaly needs to account for the periodical nature of the element, where errors vary between $[0,\pi]$.
0.D.3.3 Results
To assess the performance and generalization ability of the follower, we conducted three experiments incorporating all perturbative forces available in OrbitZoo, including harmonic gravity fields, third-body forces (Sun and Moon), solar radiation pressure, and atmospheric drag.
As detailed in Table 8, the first experiment evaluates the model under the same initial orbital elements used during training but with all forces applied. In the second experiment, the agent starts in an orbit with a larger semi-major axis, higher eccentricity and inclination, and a different initial anomaly. Finally, the third experiment tests the agent in a lower orbit (smaller semi-major axis), with reduced eccentricity and a new initial anomaly. While Experiment 1 assesses the agent’s generalization to different perturbative forces, Experiment 2 evaluates its adaptability when starting in an orbit with significantly different characteristics. Lastly, Experiment 3 tests the agent’s ability to escape Earth’s stronger gravitational potential by applying larger thrusts, which requires greater fuel consumption.
Table 8: Chase Target: Initial Keplerian elements of the follower for the different experiments.
Experiment a e i $\boldsymbol{\omega}$ $\boldsymbol{\Omega}$ M
Experiment 1 10000 + $R_{E}$ km 0.1 $5.0^{\circ}$ $10.0^{\circ}$ $10.0^{\circ}$ $10.0^{\circ}$
Experiment 2 15000 + $R_{E}$ km 0.5 $8.0^{\circ}$ $10.0^{\circ}$ $10.0^{\circ}$ $90.0^{\circ}$
Experiment 3 5000 + $R_{E}$ km 0.001 $5.0^{\circ}$ $10.0^{\circ}$ $180.0^{\circ}$ $180.0^{\circ}$
The most straightforward way to assess the agent’s performance is by tracking the immediate reward over a full episode. Since the reward represents an error, the agent’s objective is to minimize it. As shown in Figure 16, the agent effectively reduces the error in all experiments, demonstrating that it has learned a successful policy for reaching the target. In Experiment 3, the reward remains constant because the agent exhausts all available fuel, maintaining the same orbit until the episode ends. This outcome is expected, as escaping Earth’s gravitational pull requires significantly more force when starting at an altitude of 5000 km compared to 10000 km.
Figure 16: Reward (or error) per step, for each experiment. An error of 0.0 indicates that the follower contains exactly the same orbital elements as the leader, including anomaly.
To gain a deeper understanding of the agent’s policy and its impact on the reward, we examine the differences in each orbital element, as shown in Figure 17. Across all experiments, the agent prioritizes increasing the semi-major axis while keeping the remaining elements as close as possible to the target.
This behavior is particularly evident in Experiment 3, where the agent initially struggles to increase the semi-major axis because it simultaneously attempts to minimize deviations in other elements while gaining altitude. In contrast, Experiment 1 is the only case where the anomaly closely matches that of the leader, whereas in the other experiments, the agent places greater emphasis on minimizing all elements. This strategy aligns with the weight distribution in the reward function, reflecting the relative importance assigned to each orbital parameter.
This experiment demonstrates that the agent learns an effective policy capable of generalizing thrust maneuvers to initial conditions it did not encounter during training, while also adapting to realistic orbital dynamics with non-conservative forces. Future improvements to this mission could include introducing additional penalties for fuel consumption in the reward function and further exploring the limits of the agent’s strategy.
Figure 17: Absolute difference between the follower and leader orbital elements per step, for each experiment.
0.D.4 Collision Avoidance
0.D.4.1 Definition
Collision avoidance missions often rely on probabilities, particularly the probability of collision (PoC) [1] between two objects. These include a maneuverable body, referred to as the target with radius $R_{T}$, and a non-maneuverable body, referred to as the chaser with radius $R_{C}$. As the propagation of these bodies is simulated, inherent uncertainties in their current positions and velocities grow over time. These uncertainties form the basis for calculating PoC and generating CDMs.
The PoC assumes that the trajectories of the bodies follow linear motion during the very short period of possible collision, neglects uncertainties in their current velocities, and considers that position uncertainties remain constant over the short timeframe [51]. By defining the relative position ($r=r_{T}-r_{C}$) and relative velocity ($\dot{r}=\dot{r}_{T}-\dot{r}_{C}$) vectors, the exact Time of Closest Approach (TCA) within the step can be calculated as:
$\text{TCA}=-\frac{r^{T}\dot{r}}{\dot{r}^{T}\dot{r}}.$ (21)
A simplified coordinate system, known as the RTN frame, is then introduced, with the chaser positioned at the origin. The unit vectors defining the RTN frame are computed as follows:
$\hat{r}=\frac{r_{C}}{\|r_{C}\|}\quad\quad\hat{n}=\frac{r_{C}\times\dot{r}_{C}}{\|r_{C}\times\dot{r}_{C}\|}\quad\quad\hat{t}=\hat{n}\times\hat{r},$ (22)
where $\hat{r}$, $\hat{t}$ and $\hat{n}$ represent the radial, transverse, and normal directions, respectively. These unit vectors form the columns of the $3\times 3$ transformation matrix $Q=[\hat{r},\hat{t},\hat{n}]$, which is used to transform the original relative position, velocity, and position covariances of the target and chaser ($\Sigma_{T}$ and $\Sigma_{C}$) into the RTN reference frame:
$r_{\text{RTN}}=Qr\quad\quad\dot{r}_{\text{RTN}}=Q\dot{r}\quad\quad\Sigma_{\text{T,RTN}}=Q\Sigma_{T}Q^{T}\quad\quad\Sigma_{\text{C,RTN}}=Q\Sigma_{C}Q^{T}.$ (23)
Next, an orthogonal coordinate system is established based on the relative vectors:
$\hat{i}=\frac{\dot{r}_{\text{RTN}}}{\|\dot{r}_{\text{RTN}}\|}\quad\quad\hat{j}=\frac{\dot{r}_{\text{T,RTN}}\times\dot{r}_{\text{C,RTN}}}{\|\dot{r}_{\text{T,RTN}}\times\dot{r}_{\text{C,RTN}}\|}\quad\quad\hat{k}=\hat{i}\times\hat{j},$ (24)
where $\hat{i}$, $\hat{j}$ and $\hat{k}$ are the unit vectors forming the orthogonal system. Given the assumption of negligible velocity uncertainty, only the $\hat{j}$ and $\hat{k}$ axes (defining the conjunction plane) are relevant for the analysis. This simplifies the problem to two dimensions by introducing a projection matrix $C=[\hat{j},\hat{k}]$.
Under the assumption that uncertainties follow a Gaussian distribution, the distribution’s parameters are given by $\mu=Cr_{\text{RTN}}$ and $\Sigma=C(\Sigma_{\text{T,RTN}}+\Sigma_{\text{C,RTN}})C^{T}$. The PoC is then calculated as the integral over a circle centered at the origin with radius $R=R_{T}+R_{C}$:
[…]
where
$\delta=\begin{bmatrix}y\\ z\end{bmatrix}-\mu$
and $|\Sigma|$ represents the determinant of matrix $\Sigma$.
In this approach, we use a faster computation method compared to the traditional one, as outlined in [20]. A PoC value greater than $10^{-6}$ indicates a high collision risk, requiring maneuvers to address it.
0.D.4.2 Environment Setup
We consider a scenario involving two spherical and isotropic bodies: a satellite (target) – with $m_{0}=250$ kg, 50 kg of fuel, $I_{sp}=3100$ s and $R_{T}=10$ meters – and a drifter (chaser) – with $R_{C}=5$ meters. The real propagation of the bodies follows a model incorporating Newtonian gravitational attraction and atmospheric drag. However, the satellite is equipped with a system capable of simulating the propagation of both bodies (and their associated uncertainties) using a simplified model based solely on Newtonian gravitational attraction, neglecting drag effects. As expected, this simulator introduces inaccuracies since it does not fully represent the real propagation dynamics. Nonetheless, it provides valuable estimates of key parameters, including the PoC, the TCA, and the miss distance, at any given moment.
The setup of the initial conditions for each episode is critical, as the satellite must begin two days prior to the initial TCA detected by its system, with a PoC exceeding $10^{-6}$. To achieve this, the initial positions of both bodies are set at the same expected location, but with opposing velocities and small uncertainties. At the start of each episode, the states of both bodies are sampled from their respective expected positions and uncertainties, followed by propagating their trajectories backward in time. The objective is for the target to learn a strategy that effectively minimizes the PoC while maintaining a trajectory close to its nominal orbit.
In practice, both bodies are considered to start with the same Keplerian elements $(a,e,i,\omega,\Omega,M)=(2000+R_{E}\text{ km},0.01,5.0^{\circ},20.0^{\circ},20.0^{\circ},10.0^{\circ})$. However, they start with opposite velocity vectors. There is also an uncertainty of 0.1 m and 0.1 m/s when sampling the initial Cartesian position and velocity of both bodies, respectively.
At each step, the satellite simulates the dynamics of both bodies until the end of the episode, considering their current states and uncertainties. The simulation outputs the TCA, miss distance, and PoC. Each step lasts 15 minutes, with termination occurring 10 steps after the initial TCA. Thrust is applied only during the first 10 seconds of each step.
The observation $o_{t}=(s_{t},M_{t},s^{\prime}_{t},M^{\prime}_{t},f_{t},P_{t})\in\mathbb{R}^{14}$ consists of the current equinoctial elements of the satellite ($s_{t}$) and the drifter ($s^{\prime}_{t}$), both bodies mean anomaly ($M_{t}$ and $M^{\prime}_{t}$), together with the current fuel ($f_{t}$) and PoC ($P_{t}$).
As in the Hohmann maneuver mission (0.D.2), mission, apart from the polar representation of the thrust, ithe action space also contains the decision ($\delta$) to perform the given thrust: $a_{t}=(T,\theta,\phi,\delta)$. This action space is limited by $(T_{\text{max}},\theta_{\text{max}},\phi_{\text{max}},\delta_{\text{max}})=(5,\pi,2\pi,1)$.
The reward function is built in a way to not apply thrusts when there’s a small PoC, and apply thrusts that do not deviate too much from the satellite’s nominal orbit ($\hat{s}$) when it is higher than $10^{-6}$, while reducing the PoC:
$r_{t}=\begin{cases}-\alpha_{1}\mathbbm{1}_{\delta>0.5}&\text{, }P_{t-1}<10^{-6}\\ -\left(\Delta s_{t}+\alpha_{2}\mathbbm{1}_{P_{t}>10^{-6}}\right)&\text{, }P_{t-1}\geq 10^{-6}\end{cases},$ (26)
where $\Delta s_{t}=w^{T}(s_{t}-\hat{s})$, with $w^{T}$ representing a vector of weights, $P_{t-1}$ and $P_{t}$ are respectively the PoC before and after the action, and $\alpha_{1}$ and $\alpha_{2}$ are weights given to each parameter.
0.D.4.3 Results
The training hyperparameters for the agent are listed in Table 9. To assess the agent’s performance, we analyze its decision-making process based on the simulated PoC, miss distance, and TCA. Additionally, we evaluate the agent’s generalization ability by enabling all perturbative forces available in OrbitZoo, including harmonic gravity fields, third body forces (Sun and Moon), solar radiation pressure, and drag. However, the satellite simulator dynamics remain unchanged, maintaining the realistic constraint of incomplete information available to the agent.
Table 9: Collision Avoidance: Training hyperparameters.
Parameter Value
Actor learning rate 0.0001
Critic learning rate 0.001
GAE $\lambda$ 0.95
Epochs 5
Discount factor ($\gamma$) 0.95
Clip ($\epsilon$) 0.5
Initial Standard Deviation 0.2
Standard Deviation Decay Rate 5000 steps
Standard Deviation Decay Amount 0.05
Minimum Standard Deviation 0.05
Experiences for Training 256
Batch Size 64
After training, Figure 18 demonstrates that the agent applies thrust early in the episode ($\delta>0.5$) for both unrealistic and realistic dynamics, successfully reducing the simulated PoC for the remainder of the episode. However, when using realistic dynamics, the simulated PoC becomes more unstable, leading to greater uncertainty for the agent when deciding when and how to apply thrust.
When the agent is given 4 days to execute the maneuvers, as shown in Figure 19, a similar strategy emerges. However, the PoC at the initially detected TCA increases unexpectedly, indicating that the simulator’s values were insufficient to predict this behavior accurately. By analyzing the changes in orbital elements, as depicted in Figure 20, we see that the agent’s strategy involves increasing the semi-major axis by 3 km while keeping the eccentricity close to its nominal value. Although the agent has the capability to make adjustments in the cross-track direction, it correctly concludes that changing the inclination is unnecessary to minimize the collision risk. A comparable policy is observed when realistic dynamics are applied.
From these experiments, we conclude that the agent can generalize its policy to realistic dynamics when given the same amount of time (2 days) to perform the maneuvers as during training. However, when the agent is provided with 4 days to take action, the uncertainty in the positions of both bodies negatively impacts the policy, leading to an unexpected increase in the PoC at the initial TCA. Future improvements could involve training the agent with a random number of days to complete the maneuvers and incorporating more informed observations to reduce uncertainty and improve decision-making.
Figure 18: Collision Avoidance in 2 days: Agent decision to apply thrust ($\delta>0.5$), Probability of Collision (POC), Miss Distance, and Time of Closest Approach (TCA) by step.
Figure 19: Collision Avoidance in 4 days: Agent decision to apply thrust ($\delta>0.5$), Probability of Collision (POC), Miss Distance, and Time of Closest Approach (TCA) by step.
Figure 20: Evolution of the semi-major axis (a) and eccentricity components ($e_{x}$ and $e_{y}$) of the orbit by step, in the 2 days and 4 days approach.
0.D.5 GEO Constellation
0.D.5.1 Definition
Unlike Low Earth Orbit (LEO) and Medium Earth Orbit (MEO), which occupy defined altitude ranges, the Geostationary Earth Orbit (GEO) is a specific orbit located 35786 km above Earth’s equator. This orbit is significant because any object within it maintains an orbital period equal to Earth’s rotational period, making it appear stationary to observers on Earth.
There are several possible types of formation that constellations of satellites in GEO can have. Here, we focus on a mission in which all agents learn to stay at equal distances around the orbit.
0.D.5.2 Environment Setup
We define an environment with $n=4$ agents in GEO, characterized by a semi-major axis of $a_{\text{GEO}}=35786+R_{E}=42164$ km, where $R_{E}$ is Earth’s equatorial radius. Each agent is a satellite with $m_{0}=250$ kg, 50 kg of which is fuel, and equipped with a thrust with $I_{sp}=3100$ s.
Each episode contains 2000 steps with step sizes of 360 s, and the dynamics are purely based on Newtonian attraction. At the start of each episode ($t=0$), agents are initialized with random mean anomalies ($M_{0}^{i},i=\{1,2,...,n\}$), representing their positions along the orbit. The objective for the agents is to learn to distribute themselves uniformly around the orbit, maintaining angular separations of $2\pi/n$ radians (90 degrees for $n=4$).
Since GEO lies in a single orbital plane near the equator, the problem’s dimensionality can be reduced. The action space is simplified to two dimensions: $a_{t}=(T,\theta)$. This configuration allows agents to perform maneuvers affecting the semi-major axis and eccentricity components of their equinoctial elements while leaving inclination unchanged. The actions are limited to $(T_{\text{max}},\theta_{\text{max}})=(5,2\pi)$.
The observation space is also reduced. Each agent observation receives its current semi-major axis, eccentricity components, remaining fuel, and the anomalies of all agents in the environment: $o_{t}=(a_{t},e_{xt},e_{yt},f_{t},M_{t}^{1},M_{t}^{2},M_{t}^{3},M_{t}^{4})$.
To calculate the collective penalty related to the difference in anomalies ($P_{M}$), we first normalize the anomalies to be within $[0,2\pi]$, followed by:
$P_{M}=\frac{1}{C^{n}_{2}}\sum_{i=1}^{n-1}\sum_{j=i+1}^{n}\max\left(0,\frac{2\pi/n-\Delta_{ij}}{2\pi/n}\right),$ (27)
where
$\Delta_{ij}=\min\left(|M_{i}-M_{j}|,2\pi-|M_{i}-M_{j}|\right).$
The reward function of each agent penalizes differences in altitude (norm of Cartesian position $r$), magnitude of thrust ($T$), and close anomalies:
$r_{t}=-(\alpha_{1}|a_{\text{GEO}}-\|r\||+\alpha_{2}T+\alpha_{3}P_{M}),$ (28)
where $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ work as customizable weights. This reward function is extensible to an arbitrary number of agents.
0.D.5.3 Results
The agents were trained independently, without shared observations for the critic network training, with the training hyperparameters provided in Table 10. After training, we tested the generalization capability of the policy by conducting Monte Carlo simulations to analyze the trends in penalty reduction for the anomalies. These trends are shown in Figure 21, using both simple dynamics (as used during training) and realistic dynamics, which include harmonic gravity fields, third body forces, solar radiation pressure, and drag.
The results indicate that the agents effectively reduce the collective penalty (anomalies distance) in both scenarios. However, their specific ability to achieve this depends on the initial configuration of the constellation. Through the OrbitZoo interface, shown in Figure 5, it is evident that the agents are successfully learning the intended objective, albeit with some imperfections.
In conclusion, even though the agents were trained independently without any information sharing, they were able to effectively learn how to maintain equal distances from one another around the GEO orbit, even when using realistic dynamics that were not part of the training. Future improvements include implementing algorithms specifically designed for Multi-Agent Reinforcement Learning (MARL), such as MAPPO, enabling information sharing through Federated Reinforcement Learning (FRL), and scaling to larger constellations of bodies to represent more realistic scenarios.
Table 10: GEO Constellation: Training hyperparameters.
Parameter Value
Actor learning rate 0.0001
Critic learning rate 0.001
GAE $\lambda$ 0.95
Epochs 5
Discount factor ($\gamma$) 0.99
Clip ($\epsilon$) 0.1
Initial Standard Deviation 0.4
Standard Deviation Decay Rate 40000 steps
Standard Deviation Decay Amount 0.05
Minimum Standard Deviation 0.05
Experiences for Training 4096
Batch Size 64
Figure 21: Anomalies Penalty ($P_{M}$) by step for thirty Monte Carlo simulations (dashed lines) using simplified dynamics (red) and realistic dynamics (blue). Solid red and blue lines represent the average for simulations with simple and realistic dynamics, respectively.
Appendix 0.E Exploratory Data Analysis of StarLink open data
0.E.1 Residuals Analysis
In this section we present further analysis to validate the precision of the OrbitZoo propagation against ephemeris data from Starlink. Residuals between real orbital data were computed by comparing overlapping predictions from ephemeris files. The residual at each time step was calculated as the Euclidean norm of the difference between the position vectors:
$\text{Residual}(t)=\|r_{\text{ephemeris}}(t)-r_{\text{Orekit}}(t)\|.$ (29)
The analysis showed that prediction errors increase significantly after the 48-hour mark. This is due to a change in propagator by the epehemeris provider. Due to the high computational cost, it shifts to a simpler model that only accounts for the oblateness of the Earth [32].
0.E.2 Residual Scenarios
The following scenarios were evaluated to understand the discrepancies between predictions:
  • Scenario 1: Full Propagation vs. Initial Ephemeris File. Residuals were computed for the entire 1000-step propagation against the first ephemeris file.
  • Scenario 2: Second Half Propagation vs. Second Ephemeris File. Residuals were computed for the second half of the propagation (500 steps) against the second ephemeris file.
  • Scenario 3: Overlapping Region Between Consecutive Files. Residuals were computed for the overlapping period (approximately 8 hours) between two consecutive ephemeris files.
Figure 22: Comparison of predictions between the first ephemeris file, starting on November 27, 2024, at 04:50 UTC, with the subsequent nine ephemeris files to form a complete three-day prediction for the satellite with NORAD ID 44744. The residuals represent the distance between the corresponding points in each comparison.
Figure 23: Comparison of predictions between the first ephemeris file, starting on November 27, 2024, at 04:54 UTC, with the subsequent nine ephemeris files to form a complete three-day prediction for the satellite with NORAD ID 44748. The residuals represent the distance between the corresponding points in each comparison.
Figure 24: Comparison of predictions between the first ephemeris file, starting on November 27, 2024, at 04:21 UTC, with the subsequent nine ephemeris files to form a complete three-day prediction for the satellite with NORAD ID 44753. The residuals represent the distance between the corresponding points in each comparison.
Figure 25: Comparison of predictions between the first ephemeris file, starting on November 27, 2024, at 05:05 UTC, with the subsequent nine ephemeris files to form a complete three-day prediction for the satellite with NORAD ID 44921. The residuals represent the distance between the corresponding points in each comparison.
Figures 22 to 25 illustrate the residuals for these scenarios. The overlapping region provides the most reliable assessment of the propagator’s accuracy, as it is based on the most recent and accurate ephemeris data.
0.E.3 Figures and Observations
Figures 26 to 33 illustrate the satellites’ positional evolution and orbital trajectories over three days. These figures demonstrate the high accuracy of Orekit’s propagator when compared to Starlink’s ephemeris predictions, particularly in the overlapping regions.
Appendix 0.F Comparison of Orekit Propagation and Ephemerides
To validate Orekit’s propagator, physical parameters were optimized using Bayesian optimization. The parameters included:
  • Drag coefficient ($C_{D}$).
  • Reflection coefficient ($C_{R}$).
  • Satellite radius.
  • Third-body gravitational forces (enabled/disabled).
The optimization minimized the Root Mean Square Error (RMSE) between Orekit’s propagated trajectory and the ephemeris data over 1000 steps (16.6 hours). The RMSE was calculated as:
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{t=1}^{N}\|r_{\text{ephemeris}}(t)-r_{\text{Orekit}}(t)\|^{2}}.$ (30)
0.F.1 Figures of Residuals and Orbits
Figures 39 to 42 present the propagated orbits compared to the ephemeris data. Oscillatory residuals were observed in the $x$, $y$, and $z$ components, increasing over time due to unmodeled perturbative forces and numerical integration errors.
Appendix 0.G Data and residual analysis
Starlink’s ephemeris files are uploaded three times a day, every 8 hours, with each file containing three days of propagated data. This means that consecutive files have overlapping predictions, with the first 8 hours of each file being considered the most accurate. By analyzing the overlapping predictions, we can evaluate Starlink’s own propagation error.
Figures 22 through 25 compare the predictions of a three-day propagation starting on November 27th at around 5:00 AM and ending three days later. The aggregation of 10 consecutive ephemeris files allows for the residual analysis of the initial three-day prediction (from the first ephemeris file) and the subsequent predictions made in the following 9 ephemeris files.
Figures  26 through 29 show the evolution of the three positional vectors for four different satellites. Since the plotted position is derived from a combination of ephemeris files, using only the most accurate segments from each file, a discontinuity occurs at the transitions between files when one ends and another begins. As expected, the position is oscillatory in nature, reaching its maximum value when the satellite is at apogee (the farthest point from Earth along the major axis of an elliptical orbit) and its minimum value when the satellite is at perigee (the closest point to Earth along the same axis).
Figure 26: Visualization of the satellite’s x, y, and z positional evolution over three days for NORAD ID 44744. Since the plotted position is derived from a combination of ephemeris files, using only the most accurate segments from each file, a discontinuity occurs at the transitions between files when one ends, and another begins.
Figure 27: Visualization of the satellite’s x, y, and z positional evolution over three days for NORAD ID 44748. Since the plotted position is derived from a combination of ephemeris files, using only the most accurate segments from each file, a discontinuity occurs at the transitions between files when one ends, and another begins.
Figure 28: Visualization of the satellite’s x, y, and z positional evolution over three days for NORAD ID 44753. Since the plotted position is derived from a combination of ephemeris files, using only the most accurate segments from each file, a discontinuity occurs at the transitions between files when one ends, and another begins.
Figure 29: Visualization of the satellite’s x, y, and z positional evolution over three days for NORAD ID 44921. Since the plotted position is derived from a combination of ephemeris files, using only the most accurate segments from each file, a discontinuity occurs at the transitions between files when one ends, and another begins.
Figure 30: Visualization of the satellite’s orbit evolution over three days for satellite with NORAD ID 44744. To display the most accurate three-day trajectory, a combination of consecutive ephemeris files was used, selecting only the most accurate segment from each ephemeris file, roughly the first 8 hours of each file.
Figure 31: Visualization of the satellite’s orbit evolution over three days for satellite with NORAD ID 44748. To display the most accurate three-day trajectory, a combination of consecutive ephemeris files was used, selecting only the most accurate segment from each ephemeris file, roughly the first 8 hours of each file.
Figure 32: Visualization of the satellite’s orbit evolution over three days for satellite with NORAD ID 44758. To display the most accurate three-day trajectory, a combination of consecutive ephemeris files was used, selecting only the most accurate segment from each ephemeris file, roughly the first 8 hours of each file.
Figure 33: Visualization of the satellite’s orbit evolution over three days for satellite with NORAD ID 44921. To display the most accurate three-day trajectory, a combination of consecutive ephemeris files was used, selecting only the most accurate segment from each ephemeris file, roughly the first 8 hours of each file.
As mentioned earlier, the first 8 hours of these ephemeris files are the most relevant as they contain the most up-to-date information at the time of their release. To assess the accuracy of our predictions using Orekit, we must also evaluate the accuracy of Starlink’s own predictions. If Starlink’s error is significant, potentially due to unforeseen events, our propagation may also exhibit a large error as a result.
A propagation of 16.6 hours (1000 steps) was performed using Orekit, Figures 34 through 37 display the residuals of this propagation for four different satellites.
The first case shown in the figures corresponds to the residuals between the propagated data and the first ephemeris file. This case is labeled as " Orekit vs Ephemeris 1" because it uses the initial ephemeris file for comparison. The propagation was initiated with the first position of this first ephemeris file.
The second case shows the residuals between the same propagated trajectory and the second ephemeris file, labeled as "ephemeris 2." This second file contains more up-to-date information compared to the first ephemeris file, as it was released 8 hours later.
Finally, the third case examines the residuals between the two consecutive ephemeris files (ephemeris 1 and ephemeris 2). These residuals represent Starlink’s propagator’s error, as the discrepancies arise solely from Starlink’s own predictions over the overlapping time frame.
As a final remark, multiple factors beyond unknown satellite characteristics can contribute to higher RMSE values.
First, limited information on how Starlink produces its ephemeris data can lead to inconsistencies when attempting to approximate the satellites’ trajectories. Researchers, [32] suggest that Starlink may use multiple propagators and even switch among them mid-propagation, introducing further uncertainty into a single ephemeris file. Consequently, reconstructing each satellite’s real orbital trajectory from Starlink’s ephemeris data may vary across different satellites.
Second, our propagation relies exclusively on the first recorded position and velocity from each Starlink ephemeris file, under the assumption that this initial state has the least accumulated error. If that first state is inaccurate, our predicted trajectory will be skewed, causing higher RMSE values.
Lastly, Starlink’s propagation benefits from detailed space weather information. As a result, some of our trajectory predictions may happen to align more closely with real conditions than others, contributing to variability in RMSE values. This same reasoning extends to the previously mentioned satellite characteristics, as well as other forces acting on the satellite.
Figure 34: Visualization of the residuals in three different cases for NORAD ID 44748. The residuals represent the distance between the corresponding points in each comparison. Case 1 (in blue) shows the residuals between the ephemeris data from the first ephemeris file and the Orekit propagation over the entire 1000 steps. Case 2 (in yellow) shows the residuals between the second ephemeris file and the Orekit propagation over the later 491 steps. Case 3 (in green) shows the residuals between the two consecutive ephemeris files during their overlapping region within the 1000 steps. The first and second ephemeris files provide position data for different time spans, with the intersection capturing the transition between them.
Figure 35: Visualization of the residuals in three different cases for NORAD ID 44753. The residuals represent the distance between the corresponding points in each comparison. Case 1 (in blue) shows the residuals between the ephemeris data from the first ephemeris file and the Orekit propagation over the entire 1000 steps. Case 2 (in yellow) shows the residuals between the second ephemeris file and the Orekit propagation over the later 503 steps. Case 3 (in green) shows the residuals between the two consecutive ephemeris files during their overlapping region within the 1000 steps. The first and second ephemeris files provide position data for different time spans, with the intersection capturing the transition between them.
Figure 36: Visualization of the residuals in three different cases for NORAD ID 44923. The residuals represent the distance between the corresponding points in each comparison. Case 1 (in blue) shows the residuals between the ephemeris data from the first ephemeris file and the Orekit propagation over the entire 1000 steps. Case 2 (in yellow) shows the residuals between the second ephemeris file and the Orekit propagation over the later 491 steps. Case 3 (in green) shows the residuals between the two consecutive ephemeris files during their overlapping region within the 1000 steps. The first and second ephemeris files provide position data for different time spans, with the intersection capturing the transition between them.
Figure 37: Visualization of the residuals in three different cases for NORAD ID 44921. The residuals represent the distance between the corresponding points in each comparison. Case 1 (in blue) shows the residuals between the ephemeris data from the first ephemeris file and the Orekit propagation over the entire 1000 steps. Case 2 (in yellow) shows the residuals between the second ephemeris file and the Orekit propagation over the later 553 steps. Case 3 (in green) shows the residuals between the two consecutive ephemeris files during their overlapping region within the 1000 steps. The first and second ephemeris files provide position data for different time spans, with the intersection capturing the transition between them.
Figure 38: Visualization of the mean position residuals between the ephemeris data and Orekit propagation over 1000 steps for the satellites with NORAD ID 44744, 44748, 44753 and 44921.
Figure 39: Comparison of the propagated orbits and the ephemeris data, for NORAD ID 44744, illustrating the overlap between the two. The cube in the left plot indicates the approximated region that is magnified in the right plot for a closer view. The Orekit propagation is represented by a solid red line, while the ephemeris data is shown as a dashed blue line.
Figure 40: Comparison of the propagated orbits and the ephemeris data, for NORAD ID 44748, illustrating the overlap between the two. The cube in the left plot indicates the approximated region that is magnified in the right plot for a closer view. The Orekit propagation is represented by a solid red line, while the ephemeris data is shown as a dashed blue line.
Figure 41: Comparison of the propagated orbits and the ephemeris data, for NORAD ID 44753, illustrating the overlap between the two. The cube in the left plot indicates the region approximated that is magnified in the right plot for a closer view. The Orekit propagation is represented by a solid red line, while the ephemeris data is shown as a dashed blue line.
Figure 42: Comparison of the propagated orbits and the ephemeris data, for NORAD ID 44921 illustrating the overlap between the two. The cube in the left plot indicates the region approximated that is magnified in the right plot for a closer view. The Orekit propagation is represented by a solid red line, while the ephemeris data is shown as a dashed blue line.